# ClearVision: Leveraging CycleGAN and SigLIP-2 for Robust All-Weather Classification in Traffic Camera Imagery

### v1

@@ Abstract
Accurate weather classification from low-quality traffic camera imagery remains a challenging task, particularly under adverse night-time conditions. In this study, we propose a scalable framework that combines generative domain adaptation with efficient contrastive learning to enhance classification performance. Using CycleGAN-based domain translation, we improve the quality of night-time images, enabling better feature extraction by downstream models. While the baseline EVA-02 model employing CLIP-based contrastive loss achieves an overall accuracy of 96.55%, it shows a significant performance gap between day-time (97.21%) and night-time conditions (63.40%). Replacing CLIP with the more lightweight SigLIP-2 (Sigmoid contrastive loss) achieves competitive overall accuracy of 94.00%, with significant improvements in night-time performance (85.90% accuracy). The combination of Vision-SigLIP-2 + Text-SigLIP-2 + CycleGAN + Contrastive training achieves the best night-time accuracy (85.90%) across all models tested, while EVA-02 with CycleGAN maintains the highest overall accuracy (97.01%) and per-class accuracies. Our findings demonstrate the potential of combining domain adaptation and efficient contrastive learning to build practical, resource-efficient weather classification systems for intelligent transportation infrastructure.
@@ I INTRODUCTION
In response to these challenges, Vision Transformers (ViTs) have emerged as a robust alternative to traditional CNN-based models. By leveraging self-attention mechanisms, ViTs can model local and global dependencies within images [10], offering superior pattern recognition capabilities. Abdelraouf et al. [11] proposed a ViT architecture with a self-spatial attention module for rain and road-surface classification, achieving strong binary classification results. More recently, Chen et al. [12] introduced MASK-CT, a hybrid model combining masked convolutional networks with transformers to enhance generalization. Earlier approaches using conventional architectures like AlexNet and ResNet-18 [13, 3] achieved reasonable accuracy but struggled with broader weather pattern recognition and generalization across diverse camera views [14, 15]. While ViT-based models offer strong performance, they often require substantial computational resources and high-resolution training data, an issue given that real-world traffic camera footage is typically low in quality, particularly under nighttime or adverse weather conditions.
@@
In this paper, we propose a novel framework that addresses both the computational burden and poor nighttime performance limitations of existing approaches. Our solution centers on Sigmoid Loss for Language-Image Pre-training (SigLIP-2-2) [16], which uses a pairwise sigmoid loss instead of the conventional cross-entropy across batch samples. This significantly reduces computational requirements while maintaining strong performance. We integrate SigLIP-2-2 with a CycleGAN-based domain adaptation technique within a contrastive learning framework, creating a comprehensive solution specifically designed for weather classification from traffic camera imagery, as illustrated in Fig. 1.
@@
The core innovation of our approach lies in the synergistic combination of these components. SigLIP-2-2 provides efficient visual-textual representation learning, CycleGAN enhances nighttime frames through domain translation, and contrastive learning further refines the embedding space to maximize discrimination between weather conditions. Using a roadside camera dataset from the Iowa Department of Transportation, we demonstrate that our Vision-SigLIP-2-2 + Text-SigLIP-2-2 + CycleGAN + Contrastive Learning framework achieves the best nighttime performance (85.90% accuracy) across all tested models while maintaining strong overall accuracy (94.00%).
@@
We benchmark our proposed framework against established models, including EVA-02 with CLIP and a standard Vision Transformer (vit-base-patch16-224-in21k) for a comprehensive evaluation. While EVA-02 with CLIP achieves slightly higher overall accuracy (97.01%), it comes at significantly higher computational cost. EVA-02 integrates CLIP with Transformer-based architectures [17] to achieve state-of-the-art performance, but relies on computationally intensive batch-wise softmax operations and the heavy memory footprint of InfoNCE loss during training [18], making it less suitable for resource-constrained deployment scenarios.
Key Contributions: This paper makes several notable contributions to the field of weather detection using traffic camera imagery:
@@
    We propose a lightweight alternative to conventional CLIP-based models by incorporating SigLIP-2-2 for efficient language-image pretraining, significantly reducing computational overhead while maintaining competitive accuracy.
@@
    We introduce an innovative CycleGAN-based domain adaptation technique specifically designed for traffic camera imagery that enhances nighttime frames, demonstrating substantial performance improvements under low-light conditions (up to 19.05% accuracy gain).
@@
    We develop and evaluate a novel contrastive learning framework that further improves the performance of SigLIP-2-based models, particularly for challenging weather conditions and low-light scenarios.
@@
    We present comprehensive benchmarking of various model configurations on a real-world traffic camera dataset, providing valuable insights into the trade-offs between computational efficiency and accuracy for weather classification tasks.
  5. 5.
    We demonstrate that while EVA-02 with CLIP achieves the highest accuracy (97.01%), our SigLIP-2-based approaches offer a more favorable balance between performance (94.00% accuracy) and computational efficiency for practical deployments.
  6. 6.
    We achieve significant computational efficiency gains with our SigLIP-2-based approach, reducing training time by 89% and inference time by 83% compared to EVA-02, making our solution ideal for resource-constrained real-world applications.
@@ II Methods
In this section, we present our framework integrating SigLIP-2 with CycleGAN for robust weather classification from traffic camera imagery.
@@ II-A1 Task and Pipeline
  1. (i)
    Initial classification: pretrained SigLIP-2 provides the first decision.
  2. (ii)
@@
  3. (iii)
    Enhancement: CycleGAN converts night images into day-like renderings.
  4. (iv)
@@ II-A2 Input and Output
Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, $x_{i}\!\in\!\mathbb{R}^{H\times W\times 3}$, $y_{i}\!\in\!\mathcal{Y}$, and an unpaired set $\mathcal{D}_{\mathrm{unpaired}}=\{x_{j}^{\mathrm{night}},\,x_{k}^{\mathrm{day}}\}$. The network predicts
@@ II-A4 Architecture Details
Our proposed SigLIP-2 framework employs a dual-encoder architecture similar to CLIP but with important modifications for improved efficiency. The vision encoder processes images as $16\times 16$ pixel patches with 12 layers and 12 attention heads, producing 768-dimensional embeddings. A similar transformer design is used for the text encoder, processing weather condition descriptions.
The key innovation of SigLIP-2 compared to CLIP is its training objective. While CLIP uses InfoNCE loss requiring large batch sizes and significant GPU memory, SigLIP-2 employs a more efficient pairwise sigmoid-based contrastive approach, enabling equivalent performance with significantly reduced computational requirements—critical for practical deployment in transportation infrastructure.
@@
The CycleGAN component consists of two generator networks ($G_{X\rightarrow Y}$ and $G_{Y\rightarrow X}$) that learn mappings between night and day domains, and two discriminator networks ($D_{X}$ and $D_{Y}$) that distinguish between real and generated images. The framework maintains cycle-consistency to ensure that translating an image from one domain to another and back preserves the original content while modifying only domain-specific features.
@@ II-A5 Loss Functions
where $\bm{e}_{i}=h_{\bm{\phi}}(f_{\bm{\theta}}(x_{i}))$, $\tau=0.1$, and $\mathcal{P}=\{(x_{i},x_{j})\mid y_{i}=y_{j}\}\cup\{(x_{i},\tilde{x}_{i})\mid x_{i}\in\mathcal{M}\}$.
@@ II-A6 Integrated Framework
The core contribution of our work is the integration of these components into a unified framework. As illustrated in Fig. 1, our system operates through the following pipeline:
@@
    Domain Translation: Night-time images undergo enhancement via CycleGAN to their day-time equivalents while preserving weather-specific features.
@@
    Feature Extraction: Both original and transformed images are processed through SigLIP-2 to extract rich visual features.
@@
    Contrastive Alignment: The contrastive learning framework ensures consistency between original and transformed representations, improving robustness to lighting variations.
@@
    Classification: The final classification layer produces probability distributions over three weather classes.
Our novel weather-preserving loss ($\mathcal{L}_{\mathrm{weather}}$) ensures the CycleGAN transformation maintains critical weather-related visual cues, addressing both computational efficiency constraints and challenging low-light conditions of traffic camera imagery.
@@ II-B1 EVA-02 Transformer
EVA-02 builds upon the Vision Transformer architecture with multi-head self-attention for capturing spatial dependencies and position-wise feedforward networks for feature transformation. It incorporates Swish Gated Linear Unit activation [19], sub-Layer Normalization, and 2D Rotary Position Embedding. Unlike CNNs, EVA-02 processes images as sequences of patches, enabling global relationship learning without convolutional biases [17]. It achieves parameter efficiency through optimized attention mechanisms and reduced hidden layer dimensionality.
@@ II-B3 Cycle Generative Adversarial Networks
This is critical for our application, as the model must avoid introducing artificial precipitation artifacts during domain conversion. Traditional supervised approaches would require aligned day-night pairs under identical weather conditions—impractical to collect. A conceptual illustration appears in Fig. 2.
Figure 2: CycleGAN architecture for night-to-day image translation: The framework employs two generator networks ($G$ and $F$) that learn mappings between night and day domains, and two discriminator networks ($D_{X}$ and $D_{Y}$) that distinguish between real and generated images in each domain. The cycle-consistency constraint ensures that translating an image from one domain to another and back preserves the original content while modifying only domain-specific features.
@@ III Dataset
We used traffic camera imagery from CCTV installations in Ames, Iowa (Iowa DoT) spanning three weather conditions: No Precipitation, Rain, and Snow. Images were standardized to 224×224 pixels and filtered following [21, 14]. Models were adapted to address challenges typical in traffic imagery, including variable lighting and weather-induced noise [15].
@@
  • SigLIP-2: Used identical dataset split and training parameters as the CLIP variant, representing our primary contribution for efficient contrastive learning.
  • Vision Transformer: Trained on a reduced subset (2,391 images) with 60-20-20 split due to computational constraints [22].
@@ IV-A Model Performance Analysis
The baseline EVA-02 model with CLIP demonstrates strong overall performance with 96.55% accuracy, 96.80% precision, and 96.65% $F_{1}$ score. However, it shows a significant performance gap between daytime (97.21% accuracy) and nighttime conditions (63.40% accuracy), highlighting the challenge of low-light imagery classification. When enhanced with CycleGAN preprocessing, EVA-02 shows substantial improvement in night-time performance (82.45% accuracy) while maintaining strong day-time results (97.45% accuracy), resulting in the best overall performance among all models with 97.01% accuracy.
@@ IV-B Vision Transformer Performance Analysis
Vision-SigLIP-2 + text-SigLIP-2 + CycleGAN + Contrastive 90.35 94.80 85.90 90.25 95.00 85.50 90.30 94.90 85.70
@@ IV-C CycleGAN Enhancement Effects
Figure 4: Overall accuracy comparison across model configurations showing EVA-02+CycleGAN achieving the highest performance at 97.01%, followed by Vision-SigLIP-2+Text-SigLIP-2+CycleGAN+Contrastive at 94.0%, while the baseline Vision Transformer performs poorly at only 55.81%. Models with CycleGAN preprocessing consistently outperform their baseline counterparts.
@@ IV-D Qualitative Analysis of Night-to-Day Conversion
Figure 6 provides a visual example of the CycleGAN’s night-to-day conversion process. The original night-time frame (left) shows poor visibility with limited contrast and detail, while the converted day-time image (right) exhibits enhanced illumination and visibility of road features while preserving the original scene’s weather characteristics. The transformation successfully addresses key challenges identified in nighttime traffic imagery [15, 24], including uneven illumination, low contrast, and color distortion, without introducing artificial weather artifacts that could mislead the classifier. This improvement in visual clarity directly translates to better feature extraction by the classification models, as quantitatively demonstrated in our performance analysis.
Figure 6: Example of CycleGAN-based night-to-day conversion showing domain adaptation capabilities: The left image shows the original night-time frame with poor visibility, limited contrast, and color distortion—typical challenges in low-light traffic monitoring. The right image displays the CycleGAN-enhanced version with significantly improved illumination, better contrast, and clearer visualization of critical weather-related features such as road conditions and precipitation patterns. Note how the transformation maintains scene geometry and weather characteristics while enhancing only illumination-dependent features, enabling better downstream classification without introducing misleading artifacts that could affect classification accuracy.
@@ IV-E Computational Efficiency Analysis
Beyond accuracy, computational efficiency is a critical factor for practical deployment in transportation infrastructure systems. Table IV presents a comparison of the computational requirements for the evaluated models. The results highlight a substantial efficiency advantage for our SigLIP-2-based approach, which requires only 40 minutes for training compared to 6 hours for EVA-02—an 89% reduction in training time. Similarly, the inference time on our test set was reduced from approximately 3 minutes for EVA-02 to just 30 seconds for our SigLIP-2-based model, representing an 83% improvement.
TABLE IV: Computational performance comparison across model architectures. Our proposed SigLIP-2 with CycleGAN and contrastive learning framework demonstrates significant efficiency advantages over both EVA-02 and the Vision Transformer models, with substantially reduced training and inference times.
@@
Vision Transformer 6 hours 3 minutes
EVA-02 6 hours 3 minutes
SigLIP-2 + CycleGAN+Constrastive 40 minutes 30 seconds
This dramatic improvement in computational efficiency while maintaining competitive accuracy (94.00% overall accuracy vs. 97.01% for EVA-02) represents a crucial advantage for real-world applications, where deployment on edge devices or resource-constrained systems is often necessary. The reduced computational footprint also translates to lower energy consumption and infrastructure costs, making our approach more sustainable and economically viable for widespread deployment.

### v2

@@ Abstract
Adverse weather conditions challenge safe transportation, necessitating robust real-time weather detection from traffic camera imagery. We propose a novel framework combining CycleGAN-based domain adaptation with efficient contrastive learning to enhance weather classification, particularly in low-light night-time conditions. Our approach leverages the lightweight SigLIP-2 model, which employs pairwise sigmoid loss to reduce computational demands, integrated with CycleGAN to transform night-time images into day-like representations while preserving weather cues. Evaluated on an Iowa Department of Transportation dataset, the baseline EVA-02 model with CLIP achieves a per-class overall accuracy of 96.55% across three weather conditions (No Precipitation, Rain, Snow) and a day/night overall accuracy of 96.55%, but shows a significant day-night gap (97.21% day vs. 63.40% night). With CycleGAN, EVA-02 improves to 97.01% per-class accuracy and 96.85% day/night accuracy, boosting night-time performance to 82.45%. Our Vision-SigLIP-2 + Text-SigLIP-2 + CycleGAN + Contrastive configuration excels in night-time scenarios, achieving the highest night-time accuracy of 85.90%, with 94.00% per-class accuracy and 93.35% day/night accuracy. This model reduces training time by 89% (from 6 hours to 40 minutes) and inference time by 80% (from 15 seconds to 3 seconds) compared to EVA-02, ideal for resource-constrained systems. By narrowing the day-night performance gap from 33.81 to 8.90 percentage points, our framework provides a scalable, efficient solution for all-weather classification, enhancing safety using existing camera infrastructure.
@@ I INTRODUCTION
In response to these challenges, Vision Transformers (ViTs) have emerged as a robust alternative to traditional CNN-based models. By leveraging self-attention mechanisms, ViTs can model local and global dependencies within images [10], offering superior pattern recognition capabilities. Abdelraouf et al. [11] proposed a ViT architecture with a self-spatial attention module for rain and road-surface classification, achieving strong binary classification results. More recently, Chen et al. [12] introduced MASK-CT, a hybrid model combining masked convolutional networks with transformers to enhance generalization. Earlier approaches using conventional architectures like AlexNet and ResNet-18 [13, 3] achieved reasonable accuracy but struggled with broader weather pattern recognition and generalization across diverse camera views [14, 15]. While ViT-based models offer strong performance, they often require substantial computational resources and high-resolution training data, an issue given that real-world traffic camera footage is typically low in quality, particularly under night-time or adverse weather conditions.
@@
In this paper, we propose a novel framework that addresses the computational burden and poor night-time performance limitations of existing approaches. Our solution centers on Sigmoid Loss for Language-Image Pre-training (SigLIP-2) [16], which uses a pairwise sigmoid loss instead of the conventional cross-entropy across batch samples. This significantly reduces computational requirements while maintaining strong performance. We integrate SigLIP-2 with a CycleGAN-based domain adaptation technique within a contrastive learning framework, creating a comprehensive solution designed explicitly for weather classification from traffic camera imagery, as illustrated in Figure 1.
@@
The core innovation of our approach lies in the synergistic combination of these components. SigLIP-2 provides efficient visual-textual representation learning, CycleGAN enhances night-time frames through domain translation, and contrastive learning further refines the embedding space to maximize discrimination between weather conditions. Using a roadside camera dataset from the Iowa Department of Transportation, we demonstrate that our Vision-SigLIP-2 + Text-SigLIP-2 + CycleGAN + Contrastive Learning framework achieves the best night-time performance (85.90% accuracy) across all tested models while maintaining strong overall accuracy (94.00%).
@@
We benchmark our proposed framework against established models, including EVA-02 with CLIP and a standard Vision Transformer (vit-base-patch16-224-in21k) for a comprehensive evaluation. While EVA-02 with CLIP achieves slightly higher overall accuracy (97.01%), it comes at significantly higher computational cost. EVA-02 integrates CLIP with Transformer-based architectures [17] to achieve state-of-the-art performance, but relies on computationally intensive batch-wise softmax operations and the heavy memory footprint of InfoNCE loss during training [18], making it less suitable for resource-constrained deployment scenarios. Key Contributions:
@@
    Lightweight SigLIP-2 framework, replacing CLIP with pairwise sigmoid loss to cut training time by 89% and inference time by 83% while achieving 94.00% overall accuracy.
@@
    CycleGAN-based domain adaptation, enhancing night-time images with a weather-preserving loss, boosting accuracy by 18.90 percentage points (from 67.00% to 85.90%).
@@
    Contrastive learning tailored to misclassified samples from one-shot classification, improving robustness in low-light conditions.
@@
    Comprehensive baseline evaluation against EVA-02 (97.01% accuracy) and Vision Transformers, showcasing a balanced trade-off between accuracy and efficiency for transportation systems.
@@ II Methods
Our framework integrates SigLIP-2 with CycleGAN for robust weather classification from traffic camera imagery.
@@ II-A1 Task and Pipeline
  1. 1.
    Initial classification: pretrained SigLIP-2 provides first decision.
  2. 2.
@@
  3. 3.
    Enhancement: CycleGAN converts night images to day-like renderings.
  4. 4.
@@ II-A2 Input and Output
Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, $x_{i}\!\in\!\mathbb{R}^{H\times W\times 3}$, $y_{i}\!\in\!\mathcal{Y}$, and unpaired set $\mathcal{D}_{\mathrm{unpaired}}=\{x_{j}^{\mathrm{night}},\,x_{k}^{\mathrm{day}}\}$. The network predicts
@@ II-A4 Architecture Details
SigLIP-2 employs a dual-encoder architecture similar to CLIP with important modifications. The vision encoder processes images as $16\times 16$ pixel patches with 12 layers and 12 attention heads, producing 768-dimensional embeddings [19]. A similar transformer design processes weather condition descriptions.
The key innovation of SigLIP-2 is its training objective. While CLIP uses InfoNCE loss requiring large batch sizes, SigLIP-2 employs a more efficient pairwise sigmoid-based contrastive approach, enabling equivalent performance with reduced computational requirements, critical for deployment in transportation infrastructure.
@@
The CycleGAN component consists of two generator networks ($G_{X\rightarrow Y}$ and $G_{Y\rightarrow X}$) that learn mappings between night and day domains, and two discriminator networks ($D_{X}$ and $D_{Y}$). The framework maintains cycle-consistency to ensure that translation preserves the original content while modifying domain-specific features.
@@ II-A5 Loss Functions
where $\bm{e}_{i}=h_{\bm{\phi}}(f_{\bm{\theta}}(x_{i}))$, $\tau=0.1$, and $\mathcal{P}=\{(x_{i},x_{j})\mid y_{i}=y_{j}\}\cup\{(x_{i},\tilde{x}_{i})\mid x_{i}\in\mathcal{M}\}$ [18].
@@
By restricting CycleGAN-transformed pairs to misclassified images ($x_{i}\in\mathcal{M}$), we focus the contrastive loss on correcting errors from the initial one-shot classification, enhancing robustness for challenging night-time samples.
@@ II-A6 Integrated Framework
Our core contribution is integrating these components into a unified framework. As illustrated in Figure  1, our system operates through:
@@
    Domain Translation: Night-time images undergo enhancement via CycleGAN while preserving weather-specific features.
@@
    Feature Extraction: Original and transformed images are processed through SigLIP-2.
@@
    Contrastive Alignment: Contrastive learning ensures consistency between representations, improving robustness to lighting variations.
@@
    Classification: The final layer produces probability distributions over weather classes.
Our novel weather-preserving loss ($\mathcal{L}_{\mathrm{weather}}$) ensures CycleGAN transformation maintains critical weather-related visual cues, addressing both computational efficiency constraints and challenging low-light conditions.
@@ II-B1 EVA-02 Transformer
EVA-02 builds upon the Vision Transformer architecture with multi-head self-attention for capturing spatial dependencies and position-wise feedforward networks for feature transformation. It incorporates Swish Gated Linear Unit activation [20], sub-Layer Normalization, and 2D Rotary Position Embedding. Unlike CNNs, EVA-02 processes images as sequences of patches, enabling global relationship learning without convolutional biases. It achieves parameter efficiency through optimized attention mechanisms and reduced hidden layer dimensionality [17].
@@ II-B3 Cycle Generative Adversarial Networks
This is critical for our application, as the model must avoid introducing artificial precipitation artifacts during domain conversion. Traditional supervised approaches would require aligned day-night pairs under identical weather conditions—impractical to collect. A conceptual illustration appears in Figure  2.
Figure 2: CycleGAN architecture for night-to-day translation: Two generators ($G$ and $F$) learn mappings between night and day domains, while two discriminators ($D_{X}$ and $D_{Y}$) distinguish real from generated images. A cycle-consistency constraint preserves content across domain translations.
@@ III Dataset
We used traffic camera imagery from CCTV installations in Ames, Iowa, provided by the Iowa Department of Transportation (DoT), encompassing three weather conditions: No Precipitation, Rain, and Snow. The dataset comprised RGB images of varying resolutions, standardized to 224$\times$224 pixels and filtered using protocols from [22, 14] for consistency. While not publicly available, similar data can be obtained from state DoTs upon request. Models were adapted to address typical challenges in traffic imagery, including variable lighting and weather-induced noise [15].
@@
  • SigLIP-2: Trained on a reduced subset (2,391 images) with 60-20-20 split due to computational constraints [23].
  • Vision Transformer: Used identical dataset split and training parameters as the SigLIP variant, representing our primary contribution for efficient contrastive learning.
@@ IV-A Model Performance Analysis
The baseline EVA-02 model with CLIP demonstrates strong overall performance with 96.55% accuracy, 96.80% precision, and 96.65% $F_{1}$ score. However, it shows a significant performance gap between daytime (97.21% accuracy) and night-time conditions (63.40% accuracy), highlighting the challenge of low-light imagery classification. When enhanced with CycleGAN preprocessing, EVA-02 shows substantial improvement in night-time performance (82.45% accuracy) while maintaining strong day-time results (97.45% accuracy), resulting in the best overall performance among all models with 97.01% accuracy.
@@ IV-B Vision Transformer Performance Analysis
Vision-SigLIP-2 + text-SigLIP-2 + CycleGAN + Contrastive 93.35 94.80 85.90 93.65 95.10 85.50 93.45 94.90 85.70
@@ IV-C CycleGAN Enhancement Effects
Figure 4: Overall accuracy comparison across model configurations. EVA-02+CycleGAN achieves the highest accuracy (97.01%), followed by Vision-SigLIP-2+Text-SigLIP-2+CycleGAN+Contrastive (94.0%), while the baseline Vision Transformer lags at 55.81%. Models with CycleGAN preprocessing consistently outperform baselines.
@@ IV-D Qualitative Analysis of Night-to-Day Conversion
Figure  6 provides a visual example of the CycleGAN’s night-to-day conversion process. The original night-time frame (left) shows poor visibility with limited contrast and detail, while the converted day-time image (right) exhibits enhanced illumination and visibility of road features while preserving the original scene’s weather characteristics. The transformation successfully addresses key challenges identified in night-time traffic imagery [15, 25], including uneven illumination, low contrast, and color distortion, without introducing artificial weather artifacts that could mislead the classifier. This improvement in visual clarity directly translates to better feature extraction by the classification models, as quantitatively demonstrated in our performance analysis.
Figure 6: Example of CycleGAN-based night-to-day conversion demonstrating domain adaptation. Up: original night-time frame with poor visibility, low contrast, and color distortion. Down: CycleGAN-enhanced image with improved illumination and clearer visualization of road and weather features, preserving scene geometry while enhancing illumination-dependent attributes for better downstream classification
@@ IV-E Computational Efficiency Analysis
Beyond accuracy, computational efficiency is vital for deploying models in transportation systems. All experiments used an NVIDIA Tesla P100 GPU with 16 GB memory. Table IV compares computational times, showing our SigLIP-2-based model trained in 40 minutes versus 6 hours for EVA-02—an 89% reduction. Inference time dropped from 3 minutes to 30 seconds, yielding an 83% improvement.
TABLE IV: Computational performance comparison across models. The proposed SigLIP-2 with CycleGAN and contrastive learning achieves faster training and inference than both EVA-02 and Vision Transformer architectures.
@@
Vision Transformer 2.5 hours 10 seconds
EVA-02 6 hours 15 seconds
SigLIP-2+CycleGAN+Contrastive 40 minutes 3 seconds
The improved computational efficiency—while maintaining competitive accuracy (94.00% vs. 97.01% for EVA-02)—stems from architectural differences between the CLIP-based EVA-02 and our SigLIP-2-based contrastive model. This offers key advantages for real-world deployment on edge or resource-limited systems. Lower computational overhead also reduces energy use and infrastructure costs, enhancing both sustainability and scalability.